OpenStudy (anonymous):

Happy Valentine's Day ! <3 lol (: A standard number cube is rolled. Find each probability. 1. P(odd number) 2. P(prime number) 3. P(not 3) 4. P(number greater than 0) 5. P(number divisible by 5) 6. P(negative number) Helpp pleaseeeee

OpenStudy (skullpatrol):

How many total numbers are there on a "standard number cube"?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

5

OpenStudy (skullpatrol):

Count again please....

OpenStudy (anonymous):

6

OpenStudy (skullpatrol):

Correct :-D

jimthompson5910 (jim_thompson5910):

That's correct, there are 6 numbers total on a standard number cube. The numbers are: 1, 2, 3, 4, 5, 6 which of these numbers are odd?

OpenStudy (skullpatrol):

Do you know what "probability" means?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

Look at the numbers 1, 2, 3, 4, 5, 6 which of these numbers are odd?

jimthompson5910 (jim_thompson5910):

can you list them out for me?

OpenStudy (anonymous):

1 , 3 , 5 , 7 , 9 , 11 ?? do i keep going ?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no just in the list of numbers from 1 to 6 so you would list out: 1, 3, 5 those are the only odd numbers we care about at the moment

jimthompson5910 (jim_thompson5910):

So we can see that there are 3 odd numbers (1, 3, 5) out of 6 numbers total (1, 2, 3, 4, 5, 6) So the chances of you rolling an odd number is 3/6 = 1/2 To find the probability, you divide the number of ways to get what you want (in this case, an odd number) by the number of outcomes total (which is the number of ways to roll a single die)

jimthompson5910 (jim_thompson5910):

Notice how 1/2 = 0.5 = 50% so basically this is saying that there's a 50% chance of rolling an odd number

OpenStudy (anonymous):

?? i didnt get da 2nd part ^

jimthompson5910 (jim_thompson5910):

how I got 1/2 = 0.5 that part ?

OpenStudy (anonymous):

wut question r u on

jimthompson5910 (jim_thompson5910):

1. P(odd number)

jimthompson5910 (jim_thompson5910):

P(odd number) means "the probability or chances of rolling an odd number"

OpenStudy (anonymous):

ohhhh

jimthompson5910 (jim_thompson5910):

yeah it's just shorthand notation to save time

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

so do you see how I got P(odd number) = 1/2

OpenStudy (anonymous):

yess

jimthompson5910 (jim_thompson5910):

ok great so how would you find P(prime number)

OpenStudy (skullpatrol):

Do you know the definition of a "prime" number?

jimthompson5910 (jim_thompson5910):

if you're not sure, list out the numbers from 1 to 6 then count the number of prime numbers in that list

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

1, 2, 3, 4, 5, 6 which of these numbers are prime numbers? Can you list them out for me?

jimthompson5910 (jim_thompson5910):

If you forget what a prime number is, a prime number is a number where the only factors are 1 and itself ex: 17 is a prime number since only 1 and 17 are factors 12 is not a prime number because 12 = 2*6 and 12 = 3*4

OpenStudy (skullpatrol):

@partyrainbow276 Are you learning anything from this question?

OpenStudy (skullpatrol):

I asked you 3 questions: #1) How many total numbers are there on a "standard number cube"? #2) Do you know what "probability" means? #3) Do you know the definition of a "prime" number? You only answered the first question on your own that is why I'm asking.

OpenStudy (skullpatrol):

Do you want to be ABLE to do this type of question on your own?